hasan buys two kinds of cloth materials for school uniforms,shirt material that costs him RS 50 per meter adn trouser material that costs him Rs 90 per meter.for every 2 meters of trouser material he buys 3 meters of shirt material. he sells the materials at 12% and 10% profit respectively. his total sale is Rs 36,660.how much trouser material did he buy?: hasan buys two kinds of cloth materials for school uniforms,shirt material that costs him RS 50 per meter adn trouser material that costs him Rs 90 per meter.for every 2 meters of trouser material he buys 3 meters of shirt material. he sells the materials at 12% and 10% profit respectively. his total sale is Rs 36,660.how much trouser material did he buy?
Answer by stanbon(18792)

hasan buys two kinds of cloth materials for school uniforms,
shirt material that costs him RS 50 per meter
and trouser material that costs him Rs 90 per meter.
------------------------------------
for every 2 meters of trouser material he buys 3 meters of shirt material.
Quantity Equation: s = (2/3)t
-------------------
he sells the materials at 12% and 10% profit respectively.
his total sale is Rs 36,660.
Value Equation: [1.12*50]s + [1.1*90]t = 36,660
56x + 99t = 36,600
-------------------
how much trouser material did he buy?
------------------
Substitute to solve for "t":
56(2/3)t +99t = 36,600
136.3333t = 36600
t = 268.46 meters
s = 178.97 meters
==============
Cheers,
Stan H.